# SenseShift6D: Multimodal RGB-D Benchmarking for Robust 6D Pose Estimation across Environment and Sensor Variations

### v1

@@ 5 Conclusion
We introduced SenseShift6D, the first sensor-aware, multimodal benchmark for 6D object-pose estimation that physically sweeps RGB exposure, RGB gain, depth-capture mode, and ambient illumination. Extensive experiments show that state-of-the-art methods suffer substantial accuracy drops under these realistic sensor shifts. Crucially, we demonstrated that adaptive sensor control at test time can recover a large portion of this lost performance – without any model retraining – and that joint RGB + depth control outperforms unimodal strategies. These findings highlight sensor control as a promising and lightweight avenue for deploying 6D-pose systems in dynamic, real-world conditions.
@@ Appendix A RGB Augmentation Settings
Appendix B Licenses
We use open-source components under various licenses for evaluation and data generation. In particular, some scripts for photorealistic rendering are adapted from BlenderProc (GPL-3.0), and evaluation pipelines integrate components from GDRNPP (Apache-2.0), ZebraPose (MIT), and HiPose (MIT). All third-party code is used in compliance with their respective licenses, with attribution provided as required.
Our dataset, SenseShift6D, is distributed under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.
Appendix C Experimental Details
C.1 Compute Resources
@@
C.2 Model Training
@@
Table 7: Configurations for various data split.
Data split Brightness (%) Sensor configuration Exposure ($\mu$s) Gain Depth-capture mode
Train-B5Def 5 Auto (1 RGB option, 1 depth option) Auto Auto Default
Train-BADef 5, 25, 50, 75, 100 Auto (1 RGB option, 1 depth option) Auto Auto Default
Train-B50Var 50 Auto + Manual (41 RGB options, 1 depth options) Auto, 2, 4, 19, 78, 312, 1250, 5000, 10000 Auto, 0, 32, 64, 96, 128 Default
Test-B5Def 5 Auto (1 RGB option, 1 depth option) Auto Auto Default, High Accuracy, High Density, Medium Density
Test-B5Var 5 Auto + Manual (21 RGB options, 4 Depth options) Auto, 9, 39, 156, 625, 2500 Auto, 16, 48, 80, 112 Default, High Accuracy, High Density, Medium Density
Appendix D Additional Evaluation Results
Table 8 presents the performance of the Tincase in Table 4, categorized by brightness level. This suggests a noticeable difference in performance at brightness 5 compared to other brightness levels. Therefore, we evaluated SenseShift6D dataset on various data splits, beyond those described in Table 2. Each split is detailed in Table 7.
Table 8: Performance of sensor control across brightness levels for Tincase.
ZebraPose GDRNPP HiPose
Brightness Test-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var
B5 AE 33.33 64.44 95.55 37.78 40.00 93.33 10.67 64.89 71.56
Rand 9.67 15.56 36.89 13.56 19.44 42.55 10.78 18.67 69.11
Oracle 80.00 97.78 97.78 88.89 91.11 100.00 75.56 71.11 71.11
B25 AE 97.78 93.33 95.56 84.44 73.33 95.56 71.11 73.33 73.33
Rand 46.11 51.11 70.00 43.89 46.22 75.11 38.67 44.56 73.33
Oracle 100.00 97.78 100.00 91.11 95.56 97.78 77.78 73.33 73.33
B50 AE 95.56 95.56 97.78 82.22 75.56 95.56 72.44 75.11 73.78
Rand 56.11 64.67 87.00 55.33 58.00 88.34 53.00 60.89 73.56
Oracle 100.00 100.00 100.00 91.11 97.78 97.78 77.78 77.78 75.56
B75 AE 97.78 93.33 97.78 84.44 80.00 95.56 73.78 74.22 76.44
Rand 54.22 69.44 89.78 50.22 62.78 92.11 46.78 69.44 75.33
Oracle 100.00 100.00 100.00 91.11 95.56 97.78 80.00 80.00 77.78
B100 AE 97.78 97.78 97.78 86.67 80.00 95.56 73.78 74.67 74.22
Rand 54.67 66.11 84.11 53.45 57.78 87.11 52.11 61.89 74.11
Oracle 100.00 100.00 100.00 91.11 93.33 95.56 77.78 77.78 77.78
Overall AE 84.43 88.88 96.88 75.11 69.78 95.11 60.36 72.44 73.87
Rand 44.15 53.37 73.55 42.29 48.85 77.05 40.27 51.09 73.09
Oracle 96.00 99.10 99.55 90.67 94.67 97.78 77.78 76.00 75.11
D.1 Trained without Environmental Shift
We conducted an experiment under the low-light condition of brightness 5, which yielded the lowest performance in brightness-wise evaluation. To eliminate environmental shifts, we set both training and testing to brightness 5, replacing the default training condition (brightness 50). As shown in Table 9, the oracle with sensor control achieves performance comparable to or even better than training and testing conducted entirely within the same brightness 5 domain, suggesting that adaptive sensor control can mitigate domain gaps without the need for retraining in the target environment.
Table 9: Comparison of models trained without environmental shift under brightness 5.
Model Train-B5Def Train-Def
Test-B5Def AE Test-B5Def AE Test-B5Var Rand Test-B5Var Oracle
ZebraPose 26.66 33.33 9.66 80.00
GDRNPP 86.67 37.78 13.56 88.89
HiPose 35.44 10.67 10.78 75.56
D.2 Trained on Diverse Sensor and Environmental Settings
In contrast to Train-Var, we partially incorporated sensor control and environmental variability into the training process. Table 10 presents AUC performance at [0:0.1], indicating that including sensor-controlled images at a fixed brightness level (Train-B50Var) can lead to improved generalization across diverse environmental conditions, when compared to using auto exposure images from multiple brightness levels (Train-BADef). This indicates that sensor control introduces rich variations which act as implicit regularization against domain shifts. Notably, even without additional training, appropriate sensor configuration at test time contributes significantly to model performance, as shown by the oracle performance under the Train-Def setting.
Table 10: AUC performance under various brightness and sensor configurations for Tincase.
ZebraPose GDRNPP HiPose
Test-Def Test-Var Test-Def Test-Var Test-Def Test-Var
Train AE AE Rand Oracle AE AE Rand Oracle AE AE Rand Oracle
Train-Def 76.80 68.14 38.20 84.62 67.34 60.93 31.51 77.50 61.58 55.90 34.78 64.22
Train-BADef 78.73 70.09 37.31 83.96 81.54 80.38 48.83 89.04 61.82 55.36 34.32 63.73
Train-B50Var 78.91 77.00 56.16 85.58 81.17 79.69 62.03 87.48 61.82 61.54 60.47 63.00
Train-Var 80.44 79.54 61.91 86.60 79.32 78.64 64.20 85.46 61.73 61.61 60.86 63.13
Appendix E Additional Qualitative Results on Depth Variations
Figure 8 visualizes how the object region appears in the depth map under each depth capture mode for all three objects. In addition, Figure 9 shows qualitative examples for Tincase, using HiPose trained on Train-Def. Given a fixed RGB image with poor visibility, the figure compares predictions and corresponding masked depth maps between the default mode and oracle-selected modes, where the default mode fails across all examples while the oracle modes often produce better results including some successful predictions.
Figure 8: Visualization of masked object depth maps across different capture modes.
@@ Appendix E Additional Qualitative Results on Depth Variations
Figure 9: Comparison of predictions under default and oracle capture modes with a fixed RGB image. (Left) Visualized object pose on RGB images: ground truth pose in red, predicted pose in green. (Right) Masked depth maps corresponding to each capture mode.

### v2

@@ 5 Conclusion
We introduced SenseShift6D, a physically captured RGB‑D benchmark that orthogonally sweeps exposure, gain, depth‑capture mode, and illumination to isolate the impact of sensor and environmental shifts on 6D object pose estimation. Across pretrained, generalizable models (GigaPose, SAM‑6D, FoundationPose) and instance‑level models (ZebraPose, GDRNPP, HiPose), we found that factory defaults are not model‑optimal: RGB‑Only control improves accuracy by +12.3 pp or more and Depth‑Only by +10.5 pp or more, while multimodal (RGB+Depth) control achieves the best results (+19.5 pp over Baseline; +7.2 pp over RGB‑Only). A single global configuration (Oracle‑Fixed) yields only +1.3 pp on average (and can underperform Baseline), underscoring the need for adaptive, per‑scene control. Beyond mean accuracy, per‑scene control reduces object‑wise disparity from 40.0 pp to 13.6 pp and lowers scene‑level variability by 8.8 pp, indicating improved stability as well as accuracy. Crucially, these gains require no retraining, positioning test‑time sensor control as a data‑ and compute‑efficient complement to model‑centric scaling.
@@ Appendix A RGB Augmentation Settings
Appendix B Experimental Details
B.1 Computing Resources
@@
B.2 Pretrained Weights for Unseen Object Pose Estimation Models
For GigaPose, we used the pretrained weight files ‘gigaPose_v1’ (coarse) and ‘refiner-rgbd-288182519’ (refiner), following instructions from GigaPose GitHub repository.
For FoundationPose, we used the pretrained weight files ‘2023-10-28-18-33-37’ (refiner) and ‘2024-01-11-20-02-45’ (scorer), following instructions from FoundationPose GitHub repository.
For SAM-6D, we used the pretrained weight file ‘vit_h’ for SAM and the ‘sam-6d-pem-base.pth’ file for the Pose Estimation Module (PEM), following instructions from SAM-6D Github repository.
B.3 Model Training for Instance-Level Pose Estimation Models
@@
Appendix C Evaluation of the full SAM-6D pipeline
Table 8 presents SAM-6D performance based on the Top-30 segmentation results produced by its ISM.
Table 8: AUC@[0:0.1] performance of SAM-6D with its Image Segmentation Model (ISM).
Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
Spray 81.17 88.93 (+7.76) 89.88 (+8.71) 78.71 (-2.46) 90.66 (+9.49)
Pringles 61.84 78.71 (+16.87) 79.76 (+17.92) 55.85 (-5.99) 83.14 (+21.30)
Tincase 63.85 82.13 (+18.28) 82.16 (+18.31) 60.02 (-3.83) 85.42 (+21.57)
Sandwich 74.66 81.13 (+6.47) 81.62 (+6.96) 71.47 (-3.19) 82.82 (+8.16)
Mouse 49.16 66.49 (+17.33) 71.46 (+22.30) 47.96 (-1.20) 74.37 (+25.21)
Overall 66.13 79.48 (+13.35) 80.97 (+14.84) 62.80 (-3.33) 83.28 (+17.15)
Appendix D Evaluation on BOP metrics
Table 9 reports the evaluation of unseen object pose estimation models—GigaPose, SAM-6D, and FoundationPose—using the BOP metrics: VSD, MSSD, and MSPD.
Table 9: Average Reacll of VSD, MSSD, and MSPD on various sensor control methods for unseen object pose estimation models: GigaPose, SAM-6D, FoundationPose.
Model Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
GigaPose
 (Nguyen et al., 2024)
Spray 71.42 / 82.68 / 84.15 84.46 / 88.68 / 90.39 83.19 / 94.05 / 94.54 77.76 / 86.15 / 87.71 91.44 / 95.22 / 96.20
Pringles 79.30 / 68.24 / 84.29 88.88 / 83.14 / 95.33 93.55 / 87.90 / 95.10 84.03 / 75.76 / 90.00 97.72 / 96.05 / 99.43
Tincase 85.04 / 95.87 / 99.20 90.87 / 99.16 / 99.96 90.72 / 97.16 / 99.60 86.87 / 94.93 / 97.38 96.48 / 99.60 / 100.0
Sandwich 80.50 / 83.28 / 86.15 88.76 / 87.54 / 87.69 91.89 / 96.31 / 99.18 85.12 / 85.64 / 86.36 98.21 / 99.95 / 100.0
Mouse 80.26 / 84.44 / 94.34 88.41 / 91.02 / 96.44 86.94 / 91.90 / 98.34 79.34 / 79.76 / 88.34 93.90 / 96.78 / 99.61
Overall 79.31 / 82.90 / 89.63 88.28 / 89.91 / 93.96 89.26 / 93.46 / 97.35 82.62 / 84.45 / 89.96 95.55 / 97.52 / 99.05
SAM-6D
 (Lin et al., 2024a)
Spray 92.47 / 96.33 / 96.53 94.96 / 98.00 / 98.15 91.00 / 93.66 / 93.66 92.77 / 96.33 / 96.53 97.43 / 99.22 / 99.17
Pringles 94.38 / 93.08 / 98.56 97.05 / 97.76 / 100.0 95.96 / 96.57 / 98.57 94.32 / 92.43 / 96.81 99.16 / 99.62 / 100.0
Tincase 97.53 / 88.95 / 90.32 96.47 / 93.33 / 94.36 96.00 / 94.67 / 95.16 97.31 / 82.15 / 83.52 97.33 / 97.20 / 97.82
Sandwich 88.74 / 94.87 / 94.87 89.42 / 94.87 / 94.92 90.19 / 94.87 / 95.38 88.65 / 94.87 / 94.87 91.32 / 95.44 / 96.31
Mouse 78.79 / 79.49 / 85.10 82.08 / 84.41 / 89.12 84.92 / 87.06 / 91.42 79.04 / 81.56 / 86.87 88.56 / 89.85 / 93.92
Overall 90.38 / 90.54 / 93.08 92.00 / 93.68 / 95.31 91.62 / 93.37 / 94.84 90.42 / 89.47 / 91.72 94.76 / 96.27 / 97.44
FoundationPose
 (Wen et al., 2024)
Spray 89.45 / 99.22 / 99.32 92.85 / 99.76 / 99.85 94.27 / 100.0 / 100.0 89.49 / 99.32 / 99.37 97.15 / 100.0 / 100.0
Pringles 93.98 / 58.29 / 75.29 96.16 / 77.00 / 86.38 96.41 / 81.10 / 89.52 94.14 / 62.43 / 77.38 99.12 / 96.10 / 98.33
Tincase 97.36 / 44.89 / 49.11 97.60 / 65.64 / 68.09 97.62 / 79.11 / 81.73 96.96 / 43.56 / 47.33 97.75 / 89.69 / 92.09
Sandwich 94.49 / 95.08 / 96.31 97.26 / 97.59 / 98.97 95.75 / 96.26 / 98.26 95.87 / 97.38 / 97.85 97.96 / 98.41 / 99.49
Mouse 88.39 / 90.59 / 94.73 91.24 / 92.59 / 97.66 90.72 / 91.95 / 96.68 89.02 / 91.56 / 96.54 93.66 / 94.15 / 98.44
Overall 92.73 / 77.61 / 82.95 95.02 / 86.52 / 90.19 94.95 / 89.68 / 93.24 93.10 / 78.85 / 83.69 97.13 / 95.67 / 97.67
Appendix E Qualitative Analysis
E.1 Qualitative Result on Pretrained, Generalizable Pose Estimation Models
Figure 789 visualize GT and predicted poses on Baseline and Oracle sensor configuration for GigaPose, SAM-6D and FoundationPose.
@@ E.1 Qualitative Result on Pretrained, Generalizable Pose Estimation Models
Figure 7: Comparison of predictions under Baseline and Oracle for GigaPose. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
Figure 8: Comparison of predictions under Baseline and Oracle for SAM-6D. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
Figure 9: Comparison of predictions under Baseline and Oracle for FoundationPose. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
E.2 Qualitative Result on Instance-Level Pose Estimation Model
Figure 10 visualizes GT and predicted poses on Baseline and Oracle sensor configuration for HiPose.
Figure 10: Comparison of predictions under Baseline and Oracle for HiPose. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
Appendix F Evaluation on RGB-D Based Models Under Different Environments
Table 10111213 shows performance of RGB-D models—HiPose, GigaPose, SAM-6D, and FoundationPose—under different illumination conditions.
Table 10: AUC@[0:0.1] performance under diverse environments of various sensor control methods for HiPose.
Environment Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
B5
Spray 83.17 85.07 (+1.90) 87.10 (+3.93) 86.07 (+2.90) 88.63 (+5.46)
Pringles 72.40 75.45 (+3.05) 78.60 (+6.19) 77.90 (+5.50) 81.12 (+8.71)
Tincase 32.96 43.00 (+10.04) 78.80 (+45.84) 76.69 (+43.73) 81.31 (+48.36)
Sandwich 66.85 69.85 (+3.00) 80.72 (+13.87) 78.82 (+11.97) 83.18 (+16.33)
Mouse 35.29 39.88 (+4.59) 64.85 (+29.56) 62.93 (+27.63) 69.88 (+34.59)
Overall 58.13 62.65 (+4.52) 78.01 (+19.88) 76.48 (+18.35) 80.82 (+22.69)
B25
Spray 87.56 88.71 (+1.15) 88.61 (+1.05) 87.59 (+0.02) 89.63 (+2.07)
Pringles 74.12 77.17 (+3.05) 79.14 (+5.02) 77.31 (+3.19) 81.74 (+7.62)
Tincase 85.20 86.33 (+1.13) 86.91 (+1.71) 85.13 (-0.07) 87.89 (+2.69)
Sandwich 80.54 83.51 (+2.97) 83.31 (+2.77) 82.97 (+2.44) 86.51 (+5.97)
Mouse 70.41 73.17 (+2.76) 74.32 (+3.90) 71.54 (+1.12) 77.29 (+6.88)
Overall 79.57 81.78 (+2.21) 82.46 (+2.89) 80.91 (+1.34) 84.61 (+5.04)
B50
Spray 87.93 88.78 (+0.85) 89.32 (+1.39) 88.07 (+0.15) 90.07 (+2.15)
Pringles 74.24 76.57 (+2.33) 79.69 (+5.45) 76.74 (+2.50) 81.62 (+7.38)
Tincase 85.11 86.69 (+1.58) 86.91 (+1.80) 85.29 (+0.18) 88.31 (+3.20)
Sandwich 81.38 84.38 (+3.00) 85.72 (+4.33) 84.03 (+2.64) 87.23 (+5.85)
Mouse 71.93 73.17 (+1.24) 74.51 (+2.59) 71.63 (-0.29) 77.44 (+5.51)
Overall 80.12 81.92 (+1.80) 83.23 (+3.11) 81.15 (+1.03) 84.93 (+4.81)
B75
Spray 88.02 89.02 (+1.00) 89.51 (+1.49) 88.02 (+0.00) 90.61 (+2.59)
Pringles 73.31 76.02 (+2.71) 79.12 (+5.81) 76.83 (+3.52) 81.86 (+8.55)
Tincase 85.27 86.49 (+1.22) 86.80 (+1.53) 85.33 (+0.07) 88.13 (+2.87)
Sandwich 82.10 85.13 (+3.03) 84.41 (+2.31) 84.38 (+2.28) 87.10 (+5.00)
Mouse 71.73 73.27 (+1.54) 74.17 (+2.44) 71.24 (-0.49) 76.07 (+4.34)
Overall 80.09 81.99 (+1.90) 82.80 (+2.71) 81.16 (+1.07) 84.75 (+4.66)
B100
Spray 87.93 89.27 (+1.34) 89.39 (+1.46) 87.95 (+0.02) 90.41 (+2.49)
Pringles 72.90 75.69 (+2.79) 78.86 (+5.95) 76.31 (+3.40) 81.40 (+8.50)
Tincase 85.38 86.87 (+1.49) 87.09 (+1.71) 85.40 (+0.02) 88.53 (+3.16)
Sandwich 82.51 85.72 (+3.21) 86.15 (+3.64) 84.31 (+1.80) 88.10 (+5.59)
Mouse 70.24 72.59 (+2.34) 73.07 (+2.83) 70.24 (+0.00) 75.39 (+5.15)
Overall 79.79 82.03 (+2.24) 82.91 (+3.12) 80.84 (+1.05) 84.77 (+4.98)
Table 11: AUC@[0:0.1] performance under diverse environments of various sensor control methods for GigaPose.
Environment Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
B5
Spray 65.76 73.34 (+7.58) 73.39 (+7.63) 66.51 (+0.75) 78.41 (+12.65)
Pringles 19.17 44.19 (+25.02) 44.64 (+25.47) 31.33 (+12.16) 66.69 (+47.52)
Tincase 60.69 71.51 (+10.82) 65.20 (+4.51) 65.89 (+5.20) 75.07 (+14.38)
Sandwich 59.90 68.82 (+8.92) 69.46 (+9.56) 61.36 (+1.46) 80.28 (+20.38)
Mouse 48.27 60.44 (+12.17) 63.80 (+15.53) 55.59 (+7.32) 74.29 (+26.02)
Overall 50.76 63.66 (+12.90) 63.30 (+12.54) 56.14 (+5.38) 74.95 (+24.19)
B25
Spray 67.44 73.20 (+5.76) 77.78 (+10.34) 70.49 (+3.05) 81.15 (+13.71)
Pringles 15.62 49.24 (+33.62) 49.36 (+33.74) 38.17 (+22.55) 73.88 (+58.26)
Tincase 63.07 72.13 (+9.06) 69.29 (+6.22) 67.29 (+4.22) 75.69 (+12.62)
Sandwich 68.08 74.18 (+6.10) 77.72 (+9.64) 75.79 (+7.71) 84.59 (+16.51)
Mouse 61.93 69.71 (+7.78) 68.15 (+6.22) 61.95 (+0.02) 77.00 (+15.07)
Overall 55.23 67.69 (+12.46) 68.46 (+13.23) 62.74 (+7.51) 78.46 (+23.23)
B50
Spray 65.51 73.71 (+8.20) 77.20 (+11.69) 71.10 (+5.59) 83.78 (+18.27)
Pringles 24.43 45.40 (+20.97) 43.33 (+18.90) 33.50 (+9.07) 70.14 (+45.71)
Tincase 64.98 72.42 (+7.44) 70.22 (+5.24) 68.31 (+3.33) 76.82 (+11.84)
Sandwich 63.92 70.46 (+6.54) 77.51 (+13.59) 78.03 (+14.11) 84.38 (+20.46)
Mouse 56.93 68.93 (+12.00) 67.98 (+11.05) 63.68 (+6.75) 78.20 (+21.27)
Overall 55.15 66.18 (+11.03) 67.25 (+12.10) 62.92 (+7.77) 78.67 (+23.52)
B75
Spray 61.41 71.10 (+9.69) 77.37 (+15.96) 70.68 (+9.27) 81.41 (+20.00)
Pringles 22.88 47.57 (+24.69) 45.69 (+22.81) 35.12 (+12.24) 72.00 (+49.12)
Tincase 66.78 73.36 (+6.58) 71.49 (+4.71) 70.78 (+4.00) 77.36 (+10.58)
Sandwich 66.72 73.00 (+6.28) 77.79 (+11.07) 78.44 (+11.72) 84.95 (+18.23)
Mouse 55.59 68.66 (+13.07) 67.63 (+12.04) 63.61 (+8.02) 76.93 (+21.34)
Overall 54.68 66.74 (+12.06) 67.99 (+13.31) 63.73 (+9.05) 78.53 (+23.85)
B100
Spray 59.59 70.71 (+11.12) 76.37 (+16.78) 70.71 (+11.12) 81.44 (+21.85)
Pringles 22.48 47.31 (+24.83) 47.64 (+25.16) 33.12 (+10.64) 70.00 (+47.52)
Tincase 66.93 73.91 (+6.98) 70.42 (+3.49) 69.38 (+2.45) 77.13 (+10.20)
Sandwich 66.10 73.15 (+7.05) 77.95 (+11.85) 76.77 (+10.67) 85.21 (+19.11)
Mouse 60.39 69.29 (+8.90) 67.54 (+7.15) 63.12 (+2.73) 77.15 (+16.76)
Overall 55.10 66.87 (+11.77) 67.98 (+12.88) 62.62 (+7.52) 78.18 (+23.08)
Table 12: AUC@[0:0.1] performance under diverse environments of various sensor control methods for SAM-6D.
Environment Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
B5
Spray 80.41 86.83 (+6.42) 84.27 (+3.86) 82.66 (+2.25) 89.27 (+8.86)
Pringles 48.98 71.45 (+22.47) 73.10 (+24.12) 65.76 (+16.78) 80.69 (+31.71)
Tincase 55.80 74.82 (+19.02) 79.64 (+23.84) 72.60 (+16.80) 83.96 (+28.16)
Sandwich 78.36 79.00 (+0.64) 80.41 (+2.05) 78.95 (+0.59) 81.15 (+2.79)
Mouse 49.85 63.24 (+13.39) 67.61 (+17.76) 58.07 (+8.22) 72.90 (+23.05)
Overall 62.68 75.07 (+12.39) 77.01 (+14.33) 71.61 (+8.93) 81.59 (+18.91)
B25
Spray 76.00 85.90 (+9.90) 77.88 (+1.88) 82.29 (+6.29) 89.49 (+13.49)
Pringles 64.29 73.93 (+9.64) 73.52 (+9.23) 66.95 (+2.66) 81.17 (+16.88)
Tincase 74.16 78.98 (+4.82) 81.73 (+7.57) 78.22 (+4.06) 84.71 (+10.55)
Sandwich 78.10 78.74 (+0.64) 80.15 (+2.05) 78.44 (+0.34) 80.87 (+2.77)
Mouse 54.54 62.59 (+8.05) 67.12 (+12.58) 60.98 (+6.44) 72.49 (+17.95)
Overall 69.42 76.03 (+6.61) 76.08 (+6.66) 73.38 (+3.96) 81.75 (+12.33)
B50
Spray 76.63 84.29 (+7.66) 82.10 (+5.47) 83.59 (+6.96) 89.56 (+12.93)
Pringles 67.45 74.74 (+7.29) 76.86 (+9.41) 71.26 (+3.81) 81.43 (+13.98)
Tincase 78.33 81.24 (+2.91) 82.13 (+3.80) 78.42 (+0.09) 85.00 (+6.67)
Sandwich 78.38 79.18 (+0.80) 79.90 (+1.52) 78.59 (+0.21) 81.54 (+3.16)
Mouse 54.98 62.76 (+7.78) 67.39 (+12.41) 60.07 (+5.09) 72.10 (+17.12)
Overall 71.16 76.44 (+5.28) 77.68 (+6.52) 74.39 (+3.23) 81.93 (+10.77)
B75
Spray 80.83 85.88 (+5.05) 83.95 (+3.12) 84.37 (+3.54) 87.20 (+6.37)
Pringles 67.48 75.38 (+7.90) 76.83 (+9.35) 71.02 (+3.54) 80.31 (+12.83)
Tincase 75.67 80.73 (+5.06) 82.93 (+7.26) 77.87 (+2.20) 84.44 (+8.77)
Sandwich 78.38 78.77 (+0.39) 80.08 (+1.70) 78.97 (+0.59) 80.82 (+2.44)
Mouse 53.93 59.54 (+5.61) 67.66 (+13.73) 59.98 (+6.05) 71.51 (+17.58)
Overall 71.26 76.06 (+4.80) 78.29 (+7.03) 74.44 (+3.18) 80.86 (+9.60)
B100
Spray 83.59 85.20 (+1.61) 85.59 (+2.00) 84.02 (+0.43) 87.22 (+3.63)
Pringles 69.21 74.31 (+5.10) 76.05 (+6.84) 71.07 (+1.86) 80.55 (+11.34)
Tincase 77.36 79.33 (+1.97) 79.16 (+1.80) 77.49 (+0.13) 81.20 (+3.84)
Sandwich 78.74 79.28 (+0.54) 80.21 (+1.47) 79.05 (+0.31) 80.97 (+2.23)
Mouse 55.80 61.61 (+5.81) 68.22 (+12.42) 60.49 (+4.69) 72.66 (+16.86)
Overall 72.94 75.95 (+3.01) 77.84 (+4.90) 74.42 (+1.48) 80.52 (+7.58)
Table 13: AUC@[0:0.1] performance under diverse environments of various sensor control methods for FoundationPose.
Environment Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
B5
Spray 82.80 84.68 (+1.88) 85.34 (+2.54) 83.44 (+0.64) 87.17 (+4.37)
Pringles 25.07 39.45 (+14.38) 45.26 (+20.19) 32.98 (+7.91) 64.10 (+39.03)
Tincase 30.87 49.44 (+18.57) 51.80 (+20.93) 41.60 (+10.73) 65.73 (+34.86)
Sandwich 81.62 84.41 (+2.79) 83.38 (+1.76) 82.77 (+1.15) 86.28 (+4.66)
Mouse 67.88 72.80 (+4.92) 75.10 (+7.22) 72.02 (+4.14) 79.54 (+11.66)
Overall 57.65 66.16 (+8.51) 68.18 (+10.53) 62.56 (+4.91) 76.56 (+18.91)
B25
Spray 84.85 86.37 (+1.52) 87.17 (+2.32) 84.90 (+0.05) 88.46 (+3.61)
Pringles 40.24 51.57 (+11.33) 51.36 (+11.12) 44.67 (+4.43) 71.07 (+30.83)
Tincase 41.24 63.33 (+22.09) 74.76 (+33.52) 47.56 (+6.32) 84.11 (+42.87)
Sandwich 83.36 86.72 (+3.36) 84.90 (+1.54) 84.10 (+0.74) 87.97 (+4.61)
Mouse 71.39 75.37 (+3.98) 75.98 (+4.59) 72.66 (+1.27) 80.51 (+9.12)
Overall 64.22 72.67 (+8.45) 74.83 (+10.61) 66.78 (+2.56) 82.43 (+18.21)
B50
Spray 85.49 86.85 (+1.36) 88.22 (+2.73) 85.32 (-0.17) 89.59 (+4.10)
Pringles 35.67 55.31 (+19.64) 50.74 (+15.07) 45.50 (+9.83) 79.00 (+43.33)
Tincase 47.24 63.93 (+16.69) 71.60 (+24.36) 54.20 (+6.96) 82.71 (+35.47)
Sandwich 83.62 87.13 (+3.51) 85.46 (+1.84) 85.00 (+1.38) 88.56 (+4.94)
Mouse 70.71 74.39 (+3.68) 75.78 (+5.07) 72.51 (+1.80) 79.88 (+9.17)
Overall 64.54 73.52 (+8.98) 74.36 (+9.82) 68.51 (+3.97) 83.95 (+19.41)
B75
Spray 85.93 87.20 (+1.27) 88.73 (+2.80) 86.12 (+0.19) 89.83 (+3.90)
Pringles 35.19 49.95 (+14.76) 60.00 (+24.81) 54.76 (+19.57) 81.88 (+46.69)
Tincase 39.51 59.67 (+20.16) 78.89 (+39.38) 52.82 (+13.31) 86.93 (+47.42)
Sandwich 83.56 86.82 (+3.26) 85.82 (+2.26) 84.92 (+1.36) 88.92 (+5.36)
Mouse 70.98 74.78 (+3.80) 75.59 (+4.61) 72.88 (+1.90) 80.10 (+9.12)
Overall 63.03 71.68 (+8.65) 77.81 (+14.78) 70.30 (+7.27) 85.53 (+22.50)
B100
Spray 84.07 87.51 (+3.44) 88.54 (+4.47) 86.00 (+1.93) 90.00 (+5.93)
Pringles 35.14 57.31 (+22.17) 46.55 (+11.41) 48.07 (+12.93) 77.40 (+42.26)
Tincase 38.89 55.69 (+16.80) 73.11 (+34.22) 57.33 (+18.44) 85.27 (+46.38)
Sandwich 83.69 86.92 (+3.23) 86.03 (+2.34) 85.21 (+1.52) 88.82 (+5.13)
Mouse 70.17 74.37 (+4.20) 76.88 (+6.71) 72.46 (+2.29) 79.85 (+9.68)
Overall 62.39 72.36 (+9.97) 74.22 (+11.83) 69.81 (+7.42) 84.27 (+21.88)
Appendix G The Use of Large Language Models (LLMs)
In this work, editing, data processing, and visualization were polished with the assistance of ChatGPT (OpenAI, 2025) and Gemini (Google, 2025).